OpenStudy (ashsmashum1290):

Find the circumference of the circle. round your answer to the nearest centimeter. A.8 cm B.16 cm C.79 cm D.157 cm

OpenStudy (ashsmashum1290):



OpenStudy (anonymous):

\[circumferance = \pi D\] d = 5 5 * 3.14 = 16 (0 dp)